The ratio between  two numbers is 4: 3 and the difference between  them is 13.Find the number 

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Sameer
 
let numbers are x and y
 
x/y = 4/3
 
and 
x – y = 13
 
hence
 
13+ y / y= 4/3
 
39 + 3y= 4y
 
y = 39 and x= 52